Operations Report – January 12th



Crew 272 Operations Report 12-1-2023

SOL: 11

Name of person filing report: Mason Kuhn

Non-nominal systems: Oven light

Notes on non-nominal systems: No additional issues, just need a new bulb. The specifications of the light bulb type have been emailed to mission support in the Sol 10 Operations Report email chain.


Spirit rover used: No

Hours: 207.1

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 111.8

Beginning charge: 100%

Ending charge: 66%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 219.3

Beginning charge: 100%

Ending charge: 61%

Currently charging: Yes

Perseverance rover used: No

Hours: 253.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

General notes on rovers: Curiosity and Opportunity were used on EVA. They performed nominally. The chargers were draped over the front bumper upon return to the Hab. Perseverance has been moved to the outpost temporarily per Mission Support’s notice.

Summary of Hab operations: The oven light was removed & inspected for its specifications. The oven currently has no bulb in it, but the bulb cover was put back on for electrical safety reasons. 1 shower was taken today, and several pasta dishes were made throughout the sol, so water usage was slightly above average. The crew has been aware that it has plenty of water left over, but in the Martian mindset, we have still been trying our best to conserve.

WATER USE: 26.2 gallons

Water (static tank): 154.6 gallons

Static tank pipe heater (on or off): On

Static tank heater (On or off): On

Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: Suits 5, 8, 9, and 10 were used on the EVA. All suits performed nominally, and all are cleaned & currently charging. All radios have been charged to full, and they have been removed from the charging docks. The EVA crew had no communication issues.

Summary of GreenHab operations: Monitored and watered GreenHab crops. Collected relevant experimental data.

WATER USE: 8 gallons

Heater: On

Supplemental light: 4 hours

Harvest: 5g beans, 15g carrots

Summary of Science Dome operations: GreenHab Officer data collection and recording took place at the center Sci Dome table. Geologist’s samples were re-organized and are being stored on a side table of the Sci Dome.

Dual split: On (from approximately 8:00 pm to 8:00 am)

Summary of RAM operations: Tools from the RAM that are still in the lower deck of the Hab are a tape measure and a Multimeter. They will be returned to the RAM on sol 12 after the last measurements are taken for final reports.

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support: N/A

Best Regards,

Mason Kuhn

Crew 272 Engineer

Copyright © The Mars Society. All rights reserved. | Main Site